Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Please help with this sum

Expert replies
by vinitrathi1 » Sun Jun 05, 2011 3:59 pm
Q.A man can hit a target once in 4 shots. If he fires 4 shots in succession, what is the probability that he will hit his target?

1. 1
2. 1/256
3. 81/256
4. 175/256
5. 144/256
Join the discussion
Source: — Problem Solving |

by cans » Sun Jun 05, 2011 4:12 pm
prob that a man can hit = 1/4 (once in 4 shots)
and thus prob that he misses = 1 - 1/4 = 3/4
4 shots fired. let H represents hit and M represents miss
this p(Hit the target) = P(H) + P(MH) + P(MMH) + P(MMMH)
= 1/4 + (3/4)*(1/4) + (3/4)(3/4)(1/4) + (3/4)(3/4)(3/4)(1/4)
1/4 + 3/16 + 9/64 + 27/256 = 175/256

IMO D
P(H) + P(MH) + P(MMH) + P(MMMH) (we get this as if he hits first, he won't need to fire any other shot. MH - misses first and hits in the 2nd. MMH - misses first 2 shots and hits in 3rd. MMMH - misses first 3 shots and hits in the 4th.)
If my post helped you- let me know by pushing the thanks button ;)

Contact me about long distance tutoring!
[email protected]

Cans!!
Join the discussion

by Brent@GMATPrepNow » Sun Jun 05, 2011 5:31 pm
vinitrathi1 wrote:Q.A man can hit a target once in 4 shots. If he fires 4 shots in succession, what is the probability that he will hit his target?

1. 1
2. 1/256
3. 81/256
4. 175/256
5. 144/256
Another way to solve this is to say that P(hit target at least once in 4 tries) = 1 - P(hit target zero times in 4 tries)

P(hit target zero times in 4 tries)=(3/4)(3/4)(3/4)(3/4)=81/256

So, P(hit target at least once in 4 tries) = 1 - 81/256
=175/256

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by pemdas » Sun Jun 05, 2011 5:54 pm
Hi Brent, don't we violate independence of events-Hit/Miss here?
P(hit target zero times in 4 tries)=(3/4)(3/4)(3/4)(3/4)=81/256
with each successive try the probability must change, BUT you keep it absolute 3/4 (Miss) at every successive try

i ask to reveal the source
i'm not sure though - OA isn't in line with what i've obtained (in spoiler) it's close to ans. 4 but deviation of 2% (175/256=68%, my calc is 70%)
[spoiler]
i would seek 3/4*2/3*1/2*1 first try
next tries 1/4*2/3*1/2*1, 3/4*1/3*1/2*1 and 3/4*2/3*1/2*1

1/4+1/12+1/8+1/4=(6+2+3+6)/24=17/24[/spoiler]
Success doesn't come overnight!
Join the discussion

by lunarpower » Wed Jun 08, 2011 2:15 am
this is another problem on which the issue is bad wording -- "a man can hit a target once in 4 shots" is ambiguous.
pemdas is interpreting this statement as "out of exactly 4 shots, exactly 1 will hit the target". brent is reading it as "he has a 1/4 chance of hitting the target, regardless of what has already happened."

brent's interpretation is more in line with my own intuition about what the writers *meant* to say, but the question is genuinely ambiguous.

in any case, though --
i can see that some posters (pemdas in particular) don't know the source of the problem.

i'll ask the same question i've asked about a thousand times:
IF YOU DON'T KNOW THE SOURCE, WHY ARE YOU USING THE PROBLEM?

do not use problems from unknown third-party sources -- most of them are TERRIBLE, and many (if not most) of them will basically cause you to un-learn, or get confused about, the things that you're supposed to learn.
Ron has been teaching various standardized tests for 20 years.

--

Pueden hacerle preguntas a Ron en castellano
Potete chiedere domande a Ron in italiano
On peut poser des questions à Ron en français
Voit esittää kysymyksiä Ron:lle myös suomeksi

--

Quand on se sent bien dans un vêtement, tout peut arriver. Un bon vêtement, c'est un passeport pour le bonheur.

Yves Saint-Laurent

--

Learn more about ron
Join the discussion

by pemdas » Wed Jun 08, 2011 10:10 am
Ron, thanks for resolving my doubt.
>> why with so many educated gmatters, we are still having an unknown source issue?
i doubt any one with much less education/schooling than us at BTG, ever buys beans at the groceries store without the origin, expiry codes
Success doesn't come overnight!
Join the discussion

by lunarpower » Thu Jun 09, 2011 1:09 am
pemdas wrote: i doubt any one with much less education/schooling than us at BTG, ever buys beans at the groceries store without the origin, expiry codes
well done.
Ron has been teaching various standardized tests for 20 years.

--

Pueden hacerle preguntas a Ron en castellano
Potete chiedere domande a Ron in italiano
On peut poser des questions à Ron en français
Voit esittää kysymyksiä Ron:lle myös suomeksi

--

Quand on se sent bien dans un vêtement, tout peut arriver. Un bon vêtement, c'est un passeport pour le bonheur.

Yves Saint-Laurent

--

Learn more about ron
Join the discussion